OpenStudy (anonymous):

Write the following equation in slope-intercept form. 3x - 2y = 5 y = -3/2 x - 5/2 y = 3/2 x - 5/2 y = -3/2x + 5/2

jigglypuff314 (jigglypuff314):

slope intercept form would want the y to be all by itself on one sides so for 3x - 2y = 5 first subtract both sides by 3x -3x -3x so -2y = 5 - 3x then divide both sides by -2 /-2 /-2 /-2 y = (3/2)x - (5/2)

OpenStudy (anonymous):

thanks